Fenwick began refusing telnet sessions. When it was rebooted, it failed to boot properly. Upon booting into minimal mode, it was determined that the system disk was full. The OPERATOR.LOG file was deleted, and Fenwick was rebooted, the system booted normally, and telnet sessions were no loger refused.
Monthly Archives: March 2003
Newest MS Security Updates Apllied to HERMIONE
The time reinstalling Office 2000 was used to apply the latest Security Updates from Microsoft. HERMIONE was rebooted.
Excel Installed on HERMIONE
In order to support an academic project, Excel 2000 was installed on HERMIONE. This involved some extra work.
PIX Firewall Reconfigured
The PIX firewall was reconfigured to allow outside access to ports 5000 through 5350 on library.ups.edu (the Library web server). This is for a second instance of the ezproxy application.
Webboard crashed yet again. System was restarted at 7:30 3/25/03.
Hermione IIS Patched to Latest Update
Hermione’s IIS instance was patched to the latest upgrade spec to fix the most recent vulnerabilities.
Trimble Hall Network Reconfigured
Jim and Larry reconfigured the Trimble Hall network topology to simplify troubleshooting in the future. The switch stacks on the first floor now feed directly to the basement MDF, whereas before, the switch stacks were daisy chained.
Merlin2 IO System Problem
On Friday, March 14th the warning indicator on Merlin2 was flashing, and the third drive was also flashing yellow. As we were in the middle of the data move from Merlin to Merlin2, we decided to push onward and deal with the problem later.
Today, the OS went BSOD. The server was shut down, then powered back up. The server came back up without any data loss and automatically started rebuilding the array onto the previously failed disk.
Move of Merlin Data to Merlin2 Completed
The move of data from Merlin to Merlin2 is now complete. Merlin has been decomissioned from file sharing.
Trimble Hall Network Malfunction V
Jim found the source of the trouble – a wireless access point in Trimble 004. The device has been taken off the network until it can be examined.
Trimble Hall Network Malfunction IV
The trouble continues in Trimble Hall. A malfunction occurred with the same symptoms. We have isolated the problem to the switch in Trimble 007. We disconnected the TH network piece by piece while monitoring connectivity with ping. When the 007 switch was disconnected, connectivity returned. Upon reconnecting the 007 switch, the network maintained connectivity. We will have to wait until the next failure to isolate the trouble to a specific port on that switch.
Trimble Hall Network Malfunction III
The Trimble Hall problem continues with a similar malfunction this afternoon. Jim dropped the Trimble Hall uplink to 10 Mbps, which cleared the problem.
Trimble Hall Network Malfunction II
The network in Trimble Hall malfunctioned again, causing campuswide network slowdowns. It appears that rebooting any switch in Trimble will clear the problem. This time, the switch in Trimble 007 was rebooted.
Trimble Hall Network Malfunction I
A switch in Trimble Hall 107 malfunctioned and caused a traffic flood into the rest of the network. Trimble Hall was disconnected from the network for approximately 1 hour while the issue was resolved. A reboot of the switch solved the problem.
Webboard hung. To correct the problem, the machine was rebooted.